The volume of gas is 50cc at 42degrees centigrade.Pressure remaining constant.What is the temperature at which the volume of gas becomes 100cc?

Grade:11

1 Answers

jyothi
9 years ago
v1=50cc v2=100cc T1=42degrees celsius T1/V1=T2/V2 T2=T1V2/V1 =315*100/50 T2=630K =357degrees centigrade